OpenStudy (konradzuse):

evaluate each integral (cos2x) / (1+sin2x) dx

OpenStudy (anonymous):

u=1+sin2x \[du=2\cos2x dx\]

OpenStudy (anonymous):

\[1/2 \int\limits_{}^{} \frac{1}{u}du =1/2 \ln(u)+C =1/2 \ln(1+\sin2x)+C\]

myininaya (myininaya):

i think it is necessary here to insert | | instead of ( )

myininaya (myininaya):

but great job john